# Safe Navigation under State Uncertainty: Online Adaptation for Robust Control Barrier Functions

### v1

@@ IV-B Simulations
IV-C Experimental Results and Comparisons
Figure 5: Experimental comparison on tracked robot. Pose is measured at the robot’s geometric center; therefore, each circular obstacle boundary is inflated by half the robot’s length. Dashed red circles denote the true safe-set boundaries, while solid red circles indicate the inflated boundaries used in the controller synthesis. Gray tubes visualize the 95% confidence envelope of the $(x,y)$ covariance along the executed paths. (Left) (Top) R-CBF-QP with online adaptation of parameters $\bar{\gamma}_{1},\bar{\gamma}_{2}$}, the proposed method. (Left) (Middle) R-CBF-QP with tunable robustness parameters, i.e., CBF-dependent $\gamma_{1}(h),\gamma_{2}(h)$}. Tunable R-CBF-QP violated the safety constraint in the presence of time-varying state uncertainty, while its performance was similar to R-CBF-QP with online adaptation in simulation. (Left) (Bottom) Robustness parameters for the adaptive and tunable R-CBF-QPs. (Right) (Middle) R-CBF-QP without robustness parameters, i.e., $\gamma_{1}\!=\!0,\gamma_{2}\!=\!0$}. This non-robust controller violates constraints, and the robot falls from the platform and tips over at  24 sec. The experiment was terminated after this unsafe behavior. This result illustrates the necessity of robustification in practice. The CBF-QP–controlled robot encounters a deadlock at the boundary of the first obstacle, marked by the purple cross ( $\times$), since the CBF-QP only uses $v$. And, R-CBF-QP with fixed robustness parameters, i.e., $\gamma_{1}\!=\!1.4,\gamma_{2}\!=\!0.3$}, this controller is more conservative. (Right) (Bottom) The value of CBF $h$ versus time.
@@ IV-C Experimental Results and Comparisons
Our algorithms operate on a tracked GVR-Bot modified from the iRobot PackBot 510 platform. A NVIDIA Jetson AGX Orin runs our Python and C++ code under the ROS1 framework. The linear and angular body velocity commands are sent from the AGX Orin to an onboard Intel Atom, which converts these commands into individual track speeds, which are regulated through high-rate controllers.
The robot’s perception system uses an Intel RealSense depth camera operating at 30 Hz. A VectorNav VN-100 provides inertial measurements. Our state estimation employs OpenVINS [24], a Visual-Inertial Odometry (VIO) that fuses IMU signals and sparse visual features via an extended Kalman filter. OpenVINS uses first-estimate Jacobians to maintain observability consistency, propagating and updating the full state covariance. The covariance estimates published by OpenVINS provide us a runtime uncertainty bound with high confidence.
We evaluated the performance of four controllers in the scenario shown in Fig. 5, in which the safe set, reference path, and obstacle geometry are the same as in simulation. The state uncertainty bound, provided by VIO, now is time–varying. The hardware study used the same hyperparameters as in Sec. IV-B, except $\alpha\!=\!1$}. Our online–adapted R-CBF-QP with $\bar{\gamma}_{1},\bar{\gamma}_{2}$}, consistently preserved safety while staying closest to the reference path. The fixed–gain R-CBF-QP exhibited the expected conservatism (earlier braking and a larger lateral tracking error), while the tunable R-CBF-QP became unsafe due to the state estimation uncertainty. The R-CBF-QP without robustness parameters, i.e., $\gamma_{1}\!=\!0,\gamma_{2}\!=\!0$} also violated safety and was terminated earlier, see Fig. 5. Quantitative outcomes are presented in Table II: our online adaptive R-CBF-QP achieves the lowest $J_{t}$ among safe controllers.
@@ V Conclusion
To consider the impact of state errors, we pre-compute tightened safety constraints based on the tube-MPC concept [25]. Specifically, we enlarge each circular obstacle by the maximum expected position error, while the edges of the rectangular boundary are shifted inward by the maximum anticipated error in the corresponding direction. The resulting set is referred to as the tightened set, that is, the 0-superlevel set of $\bar{h}_{i}(\mathbf{x})$}, which is defined as the tightened version of our CBFs $h_{i}(\mathbf{x})$} for $i\!\in\!\{1,\ldots,6\}$}. Therefore, when ${\bar{h}_{i}(\mathbf{x})}\geq 0$, the robot remains within the safe set, even in the presence of estimation errors. Then, we define an R-COCP:

### v2

@@ IV-C Experimental Results and Comparisons
Our algorithms operate on a tracked GVR-Bot modified from the iRobot PackBot 510 platform. An NVIDIA Jetson AGX Orin runs our Python and C++ code under the ROS1 framework. The linear and angular body velocity commands are sent from the AGX Orin to an onboard Intel Atom, which converts these commands into individual track speeds, which are regulated through high-rate controllers.
The robot’s perception system uses an Intel RealSense depth camera operating at 30 Hz. A VectorNav VN-100 provides inertial measurements. Our state estimation employs OpenVINS [24], a visual-inertial odometry (VIO) that fuses IMU signals and sparse visual features via an extended Kalman filter. OpenVINS uses first-estimate Jacobians to maintain observability consistency, propagating and updating the full state covariance. The covariance estimates published by OpenVINS provide us with a runtime uncertainty bound. We use a 95% confidence interval.
We evaluated the performance of four controllers in the scenario shown in Fig. 5, in which the safe set, reference path, and obstacle geometry are the same as in simulation. The state uncertainty bound, provided by VIO, is now time–varying. The hardware study used the same hyperparameters as in Section IV-B except that we employed a coarser $80\times 80$ grid for the tunable parameters and set $\alpha=1$. On our platform, solving the R-CBF-QP with this grid takes approximately $13~\mathrm{ms}$ per control step, within the $20~\mathrm{ms}$ budget of the $50~\mathrm{Hz}$ control loop. Our online–adapted R-CBF-QP with $\bar{\gamma}_{1},\bar{\gamma}_{2}$}, consistently preserved safety while staying closest to the reference path. The fixed–gain R-CBF-QP exhibited the expected conservatism (earlier braking and a larger lateral tracking error), while the tunable R-CBF-QP became unsafe in hardware under time-varying state estimation uncertainty. Note that the tunable R-CBF-QP uses the CBF-dependent gains $\gamma_{1}(h),\gamma_{2}(h)$ given in (19), which do not adapt to the time-varying $\delta(\hat{\mathbf{x}}(t))$; this can leave insufficient robustness on hardware and lead to safety violations. The R-CBF-QP without robustness parameters, i.e., $\gamma_{1}\!=\!0,\gamma_{2}\!=\!0$} also violated safety and was terminated earlier, see Fig. 5. Quantitative outcomes are presented in Table II: our online adaptive R-CBF-QP achieves the lowest $J_{t}$ among safe controllers.
@@ V Conclusion
To consider the impact of state errors, we pre-compute tightened safety constraints based on a tube model predictive control (MPC) concept [25]. Specifically, we enlarge each circular obstacle by the maximum expected position error, while the edges of the rectangular boundary are shifted inward by the maximum anticipated error in the corresponding direction. The resulting set is referred to as the tightened set, that is, the 0-superlevel set of $\bar{h}_{i}(\mathbf{x})$}, which is defined as the tightened version of our CBFs $h_{i}(\mathbf{x})$} for $i\!\in\!\{1,\ldots,6\}$}. Therefore, when ${\bar{h}_{i}(\mathbf{x})}\geq 0$, the robot remains within the safe set, even in the presence of estimation errors. Then, we define a robust constrained optimal control problem (R-COCP):